OpenStudy (igigighjkl):

Jade used Fraction 3 over 4 yard of fabric to make a scarf. Can she make 2 of these scarves with Fraction 1 and 4 over 5 yards of fabric, and why? No, because the quotient of Fraction 3 over 4 ÷ Fraction 1 and 4 over 5 is 1 and 7 over 20 No, because the quotient of Fraction 1 and 4 over 5 ÷ Fraction 3 over 4 is 1 and 7 over 20 Yes, because the quotient of Fraction 3 over 4 ÷ Fraction 1 and 4 over 5 is 2 and 2 over 5 Yes, because the quotient of Fraction 1 and 4 over 5 ÷ Fraction 3 over 4 is 2 and 2 over 5 Note: I only want HELP!!!!

OpenStudy (igigighjkl):

@pooja195

OpenStudy (igigighjkl):

@pooja195 please HELP me only.

OpenStudy (igigighjkl):

@ShiBuringa help me please!!!

OpenStudy (shiburinga):

Hm...

OpenStudy (igigighjkl):

@pooja195

OpenStudy (shiburinga):

What do you think it is?

OpenStudy (anonymous):

Let's examine the problem. :D The amount of fabric used to make ONE scarf is: \[\frac{ 3 }{ 4 }\]

OpenStudy (igigighjkl):

Don't Know at all not a clue hopeless got noth' in

OpenStudy (anonymous):

Oh, and the units are yards. :)

OpenStudy (shiburinga):

@Michele_Laino

OpenStudy (igigighjkl):

@pooja195

OpenStudy (shiburinga):

Thanks @Michele_Laino

OpenStudy (shiburinga):

bye gtg :P

OpenStudy (igigighjkl):

Thanks Guys I'll just do it myself. o.o

OpenStudy (michele_laino):

hint: in order to make 2 scarves we need of \(2 \cdot 3/4=3/2\) of yard, and we have \(\Large 1\frac{4}{5} = 1 + \frac{4}{5} = \frac{9}{5}\) of yard

OpenStudy (michele_laino):

so, we have to compare these two fractions: \(3/2\) and \(9/5\)

OpenStudy (anonymous):

Alright, I've got it. :D First, let's examine the problem: It takes 3/4 yards of fabric to make one scarf. Therefore, to make TWO scarves, it will take twice the amount of fabric. So: \[2 \times \frac{ 3 }{ 4 } = \frac{ 6 }{ 4 }\] We can reduce this fraction to: \[\frac{ 3 }{ 2 }\] So, it takes at least 3/2 yards of fabric to make TWO scarves. In the problem, Jade wants to use the following amount to make two scarves: \[1 \frac{ 4 }{ 5}\] If that number is equal to or greater than 3/2, then Jade will have enough fabric to make two scarves. Converting the mixed number into an improper fraction, we get: \[\frac{ 9 }{ 5 }\] Now, we must figure out if this number is equal to or greater than 3/2. We can do that by comparing them. First, we must get these two fractions to have the same denominator. We can multiply 3/2 by 5: \[\frac{ 3 }{2 } \times 5 = \frac{ 15 }{ 10 }\] We can also multiply 9/5 by 2: \[\frac{ 9 }{ 5 } \times 2 = \frac{ 18 }{ 10 }\] Now that the fractions have the same denominator, we can clearly see which is bigger: \[\frac{ 15 }{ 10 } < \frac{ 18 }{ 10}\] 18/10, or 9/5, is BIGGER. The amount of fabric that Jade has will be MORE THAN ENOUGH to make the two scarves. Therefore, YES, she will be able to make them. However, you see that TWO answer choices have ''yes'' in them. They are C and D. The second part of C states: \[\frac{ 3 }{ 4}\div \frac{ 9 }{ 5 }= 2 \frac{ 2 }{ 5 }\] The second part of D states: \[\frac{ 9 }{ 5 } \div \frac{ 3 }{ 4 }= 2 \frac{ 2 }{ 5 }\] Only one of these statements can be true. Let's try answer choice D. We can use the reciprocal of 3/4 to change this division problem into a multiplication problem: \[\frac{ 9 }{ 5 } \times \frac{ 4 }{ 3 } = \frac{ 36 }{ 15 }\] We can reduce this fraction to \[\frac{ 12 }{ 5 }\] Turning this into a mixed number gets us: \[2\frac{ 2 }{ 5 }\] The statement in answer choice D is correct. Therefore, the correct answer choice is D. If you need more help, please let me know. :D